Sandpiper Company has 15,000 shares of cumulative preferred 3% stock, $50 par and 50,000 shares of $10 par common stock. The following amounts were distributed as dividends:Sandpiper Company has 15,000 shares of cumulative preferred 3% stock, $50 par and 50,000 shares of $10 par common stock. The
Answer #1

The preferred shares are cumulative it means that if the dividend is not paid for the year then it will be carried forward and paid in next year and the remaining amount of dividend is distributed to common stock holders.

Here 15000 cumulative preferred shares are there, dividend is 3% on the par value of $50.

Thus dividend payable on per share = Rate * par value * Number of shares = 15000 * 50 * 3% = $ 22,500

Thus, for Year-1.

Dividend Decared and distruibuted = $ 33800, since it is more than $ 22500 so the preferred shareholders will receive the full dividend of $ 22500 and the remaining of $ 11300 ( 33800 - 22500 ) will be distributed to common stockholders.

Thus Preferred stocks dividend per share = Dividend distributed / Number of share = $ 22500 / 15000 = $ 1.50.

Common stock dividend per share = $ 11300 / 50000 = 0.23

Thus, for Year-2

Dividend Decared and distruibuted = $ 18000, since it is less than $ 22500 so the preferred shareholders will receive the full dividend of $ 18000 and no dividend will be paid to the common stock holders , further the preferred shares are cumulative so short paid dividend of $ 4500 ( 22500 - 18000 ) will be paid in next year first.

Thus Preferred stocks dividend per share = Dividend distributed / Number of share = $ 18000 / 15000 = $ 1.20

Common stock dividend per share = 0

Thus, for Year-3

Dividend Decared and distruibuted = $ 67500 since it is more than $ 22500 so the preferred shareholders will receive the full dividend of $ 22500 and the opening cumulative balance of $ 4500 of preferred stocks dividend will also be received , so total dividend paid to preferred stocks = $ 22500 + 4500 = $ 27000

The remaining of $ 40500 ( 67500 - 22500 - 4500 ) will be distributed to common stockholders.

Thus Preferred stocks dividend per share = Dividend distributed / Number of share = $ 27000 / 15000 = $ 1.80.

Common stock dividend per share = $ 40500 / 50000 = 0.81.

Year-1 Year-2 Year-3
Preferred stocks (Dividend per share) 1.50 1.20 1.80
Common stocks (Dividend per share) 0.23 0 0.81

Feel free to comment doubts if any..